# JWST COMPASS: Insights into the Systematic Noise Properties of NIRSpec/G395H From a Uniform Reanalysis of Seven Transmission Spectra

### v1

@@ 6.2 Future Observations
Sub-Neptunes with detections of molecules in their atmospheres tend to sit on the upper edge of the mass-radius diagram at large radii relative to their masses. Detections of transmission features for sub-Neptunes with less inflated radii have been less forthcoming, with the possible exception of K2-18 b. Transmission features have been elusive for super-Earths, with only L 98-59 b showing tentative features (Bello-Arufe2025). This suggests the possibility that observers have been overly optimistic when estimating the number of transits necessary to detect atmospheres of all but the most inflated small planets. This may be due to a more frequent occurrence of low pressure aerosols and high metallicity atmospheres than was expected prior to the launch of JWST, the impact of stellar variability and surface inhomogeneities as well as other sources of correlated noise in JWST timeseries, and systematic overprediction of the transit precision from tools like pandexo. If we wish to begin building a picture of atmospheric composition across the mass-radius diagram, we will need to observe additional transits of those planets that have already been studied, and to allocate greater numbers of transits to new targets.
@@
We investigate the prospects for detecting transmission features for our sample of seven COMPASS planets with additional observations using either NIRSpec/G395H or NIRISS/SOSS, under the assumption that these worlds possess clear, high metallicity atmospheres. We consider adding an additional transit observations to the existing transits observed for the COMPASS program. Assuming a 300 Solar metallicity, C/O=1 Solar clear atmosphere and transit observations for each planet, we estimate , the rejection threshold for a flat model. In order to roughly marginalize over realizations of the white noise in the spectrum, we simulate 10,000 spectra with independent white noise realizations to obtain a 95% confidence interval for . For the NIRSpec simulations we compute the uncertainty on the spectrum by scaling the measured uncertainty for the transits by a factor of . For the NIRISS simulations we use pandexo to determine the expected precision of the spectrum. We then inflate the precision by a factor of 1.2 to account for the fact that pandexo underestimates transit depth uncertainties by assuming a flat-bottomed transit rather than a limb-darkened transit. The factor of 1.2 is chosen to match the observed discrepancy between the pandexo precision and the real precision for NIRSpec, according to Espinoza2023.
@@
As discussed in Section 5 and illustrated in Figure 21, a picture is beginning to emerge of the diversity of sub-Neptune atmospheres. In particular, the planets GJ 9827 b and TOI-421 b both have spectra dominated by water absorption, with smaller relative abundances of CH4 and CO2 than for TOI-270 d and K2-18 b. In order to understand how this diversity impacts prospects for detecting atmospheres of these seven COMPASS targets, we carry out the same simulations described above for atmospheric compositions that are 50% steam/50% H2. The results of these simulations are shown in Figure 24. For these atmospheres NIRISS/SOSS and NIRSpec/G395H perform similarly, with the exception of TOI-836 c and TOI-836 b for which NIRISS requires approximately two fewer transits to detect the steam atmosphere than does NIRSpec. The improved performance of NIRISS over NIRSpec for the steam atmospheres over the equilibrium atmospheres in Figure 23 is a result of the fact that NIRISS captures several prominent water features at wavelengths shorter than 2.8 m which are missed by NIRSpec.
@@
As discussed in previous COMPASS publications and illustrated in Figure 14, the degeneracy between metallicity (or mean molecular weight) and opaque pressure level presents a challenge for interpreting flat transmission spectra. Potential means of breaking this degeneracy include atmospheric escape observations, which can indicate the presence of an atmosphere and provide some constraints on metallicity, emission spectroscopy through secondary eclipse observations, which can also be used to demonstrate the presence of an atmosphere, and additional transit observations to obtain more precise transmission spectroscopy. Emission spectroscopy is likely to be challenging for most off these small, temperate planets. The emission spectroscopy metric (ESM) put forward by Kempton2018, which is a measure of the expected signal-to-noise ratio of the planet’s secondary eclipse, ranges between 1.6 for TOI-776 c and 13.8 for L 168-9 b. For reference, we compute the ESM for GJ 1214 b, which has been successfully observed in emission with a single transit observed by MIRI LRS (Kempton2024), to be 26.4 using the planetary and stellar parameters from Mahajan2024. These low ESM values indicate that, while the hottest planets in our sample may permit emission spectroscopy in a manageable number of transits, transmission spectroscopy will be an essential tool if we wish to break the opaque pressure level/metallicity degeneracy and measure the atmospheric compositions of these planets.
@@
It should be noted that none of these simulations account for the possibility that NIRISS/SOSS observations could be impacted by stellar contamination via the transit light source effect (Rackham2018), which can alter the transmission spectrum at short wavelengths and potentially mimic a water vapor signature (see, e.g. Moran2023). Another factor for NIRISS/SOSS observations not accounted for in our analysis is the possibility of a detection of the Rayleigh scattering slope of haze particulates in the atmosphere. Though potentially degenerate with both water features and stellar contamination, a successful detection of a Rayleigh scattering slope could confirm the presence of a haze, providing an additional path towards breaking the opaque pressure level/metallicity degeneracy.
@@
7 Conclusions
We have carried out independent reductions of the first seven targets from the COMPASS program using the ExoTiC-JEDI code. We introduce a systematics model that accounts for variations in the trace morphology via the principal components of the relative pixel flux timeseries. We demonstrate that this systematics model reduces red noise in the residuals of the white light curves and improves the Akaike and Bayesian information criteria, with the improvement being particularly profound for the observations with three groups per integration. An analysis of the posteriors from the white light curve fitting confirms that the systematic effects are largest for the three-group observations. We also observe that the NRS1 detector suffers from systematics to a greater degree than NRS2.
@@ 7 Conclusions
We derived new limits on metallicity and opaque pressure level assuming thermochemical equilibrium for each of the seven targets. We showed that the lower limits on metallicity are in broad agreement with those obtained in earlier studies with a few exceptions as noted in Section 5.
@@
By comparing best-fit models from the set of small planets with JWST observations of transmission features in the near-infrared to hypothetical chemically-identical atmospheres around each of the COMPASS targets, we demonstrate that only the lowest mean molecular weight sub-Neptune or super-Earth atmospheres known, those of TOI-421 b and K2-18 b, would be detectable around all of the planets in the sample. In other words, we confirm that none of the planets in this sample have atmospheres that are analogs to K2-18 b or TOI-421 b. Notably this includes TOI-836 c, which is a close analog of TOI-421 b in mass and radius, though it is substantially cooler with an equilibrium temperature of compared to TOI-421 b’s . The atmosphere of TOI-421 b is ruled out for TOI-836 c at . If these planets have similar atmospheric chemistry, it would therefore imply the presence of high-altitude aerosols in TOI-836 c’s atmosphere similar to those observed for GJ 1214 b. With two to five transits of TOI-836 c observed with NIRISS/SOSS, a 50% steam/50% H2 atmosphere could be ruled out, while transits would be required to rule out a 100 Solar metallicity atmosphere with an opaque pressure level bar with NIRSpec/G395H. Using the same methods described in Section 6.2 we find that an atmosphere analogous to that of GJ-1214 b for TOI-836 c would be undetectable with fewer than several hundred NIRSpec/G395H transits due to its smaller transmission spectroscopy metric (Kempton2018). As a result, breaking the aerosol-metallicity degeneracy by detecting aerosols in the astmosphere of TOI-836 c would likely require the use of NIRISS/SOSS to detect the Rayleigh scattering slope at wavelengths shorter than those accessible to G395H.
The heavier atmospheres of TOI-270 d, L 98-59 b, and GJ 9827 b can all be ruled out for some of the seven COMPASS targets but are compatible with others, while the very high mean molecular weight, hazy atmosphere of GJ 1214 b is compatible with all seven of our spectra. In particular we note that TOI-270 d’s atmospheric composition is strongly ruled out for four of the seven targets (), weakly ruled out for TOI-836 c and TOI-776 b () and consistent only with TOI-776 c. Teske2025 also finds that TOI-776 c’s spectrum is compatible with TOI-270 d’s atmospheric composition, indicating the need for additional observations to rule out the hypothesis that these two planets, which have similar masses, radii, and equilibrium temperatures, have analogous atmospheres.
@@
In addition to the software cited in the text, we acknowledge the use of the following software in this work: astropy (astropy), scipy (scipy), numpy (numpy), emcee (emcee), matplotlib (matplotlib), scikit-learn (scikit-learn), batman (Kreidberg2015).

### v2

@@ VI.2 Future Observations
Sub-Neptunes with detections of molecules in their atmospheres tend to sit on the upper edge of the mass-radius diagram at large radii relative to their masses. Detections of transmission features for sub-Neptunes with less inflated radii have been less forthcoming, with the possible exception of K2-18 b. Transmission features have been elusive for super-Earths, with only L 98-59 b showing tentative features (Bello-Arufe et al., 2025). This suggests the possibility that observers have been overly optimistic when estimating the number of transits necessary to detect atmospheres of all but the most inflated small planets. This may be due to a more frequent occurrence of low pressure aerosols and high metallicity atmospheres than was expected prior to the launch of JWST, the impact of stellar variability and surface inhomogeneities as well as other sources of correlated noise in JWST timeseries, and systematic overprediction of the transit precision from tools like pandexo. If we wish to begin building a picture of atmospheric composition across the mass-radius diagram, we will need to observe additional transits of those planets that have already been studied, and to allocate greater numbers of transits to new targets.
@@
We investigate the prospects for detecting transmission features for our sample of seven COMPASS planets with additional observations using either NIRSpec/G395H or NIRISS/SOSS, under the assumption that these worlds possess clear, high metallicity atmospheres. We consider adding an additional transit observations to the existing transits observed for the COMPASS program. Assuming a 300 Solar metallicity, C/O=1 Solar clear atmosphere and transit observations for each planet, we estimate , the rejection threshold for a flat model assuming no detector offsets for the NIRSpec/G395H observations. In order to roughly marginalize over realizations of the white noise in the spectrum, we simulate 10,000 spectra with independent white noise realizations to obtain a 95% confidence interval for . For the NIRSpec simulations we compute the uncertainty on the spectrum by scaling the measured uncertainty for the transits by a factor of . For the NIRISS simulations we use pandexo to determine the expected precision of the spectrum. We then inflate the precision by a factor of 1.2 to account for the fact that pandexo underestimates transit depth uncertainties by assuming a flat-bottomed transit rather than a limb-darkened transit. The factor of 1.2 is chosen to match the observed discrepancy between the pandexo precision and the real precision for NIRSpec, according to Espinoza et al. (2023).
@@
As discussed in Section V and illustrated in Figure 25, a picture is beginning to emerge of the diversity of sub-Neptune atmospheres. In particular, the planets GJ 9827 d and TOI-421 b both have spectra dominated by water absorption, with smaller relative abundances of CH4 and CO2 than for TOI-270 d and K2-18 b. In order to understand how this diversity impacts prospects for detecting atmospheres of these seven COMPASS targets, we carry out the same simulations described above for atmospheric compositions that are 50% steam/50% H2. The results of these simulations are shown in Figure 28. For these atmospheres NIRISS/SOSS and NIRSpec/G395H perform similarly, with the exception of TOI-836 c and TOI-836 b for which NIRISS requires approximately two fewer transits to detect the steam atmosphere than does NIRSpec. The improved performance of NIRISS over NIRSpec for the steam atmospheres over the equilibrium atmospheres in Figure 27 is a result of the fact that NIRISS captures several prominent water features at wavelengths shorter than 2.8 m which are missed by NIRSpec.
An additional complicating factor for NIRSpec/G395H observations of steam atmospheres is the degeneracy between the slope of the m water feature extending across the NRS1 portion of the spectrum and the slopes that can arise from stellar contamination in this same wavelength range. This effect may further advantage NIRISS/SOSS over NIRSpec/G395H when it comes to steam atmospheres, particularly for M dwarfs where stellar contamination is expected to be a limiting factor. The presence of the detector gap and the associated offset between the NRS1 and NRS2 portions of the spectrum may also present a difficulty for constraining the presence of water vapor in an atmosphere: Because this gap overlaps with the edge of the water feature slope, it is possible for a degeneracy to arise between the amplitude of the water feature and the detector offset (see, e.g. Coulombe et al., 2025).
@@
As discussed in previous COMPASS publications and illustrated in Figure 17, the degeneracy between metallicity (or mean molecular weight) and opaque pressure level presents a challenge for interpreting flat transmission spectra. Potential means of breaking this degeneracy include atmospheric escape observations, which can indicate the presence of an atmosphere and provide some constraints on metallicity (e.g. Zhang et al., 2025), emission spectroscopy through secondary eclipse observations or phase curves, which can also be used to demonstrate the presence of an atmosphere (Kempton et al., 2023), and additional transit observations to obtain more precise transmission spectroscopy. Emission spectroscopy is likely to be challenging for most off these small, temperate planets. The emission spectroscopy metric (ESM) put forward by Kempton et al. (2018), which is a measure of the expected signal-to-noise ratio of the planet’s secondary eclipse, ranges between 1.6 for TOI-776 c and 13.8 for L 168-9 b. For reference, we compute the ESM for GJ 1214 b, which has been successfully observed in emission with a single transit observed by MIRI LRS (Kempton et al., 2023), to be 26.4 using the planetary and stellar parameters from Mahajan et al. (2024). These low ESM values indicate that, while the hottest planets in our sample may permit emission spectroscopy in a manageable number of transits, transmission spectroscopy will be an essential tool if we wish to break the opaque pressure level/metallicity degeneracy and measure the atmospheric compositions of these planets.
@@
It should be noted that none of these simulations account for the possibility that NIRISS/SOSS observations could be impacted by stellar contamination via the transit light source effect (Rackham et al., 2018), which can alter the transmission spectrum at short wavelengths and potentially mimic a water vapor signature (see, e.g. Moran et al., 2023). Another factor for NIRISS/SOSS observations not accounted for in our analysis is the possibility of a detection of the Rayleigh scattering slope of haze particulates in the atmosphere. Though potentially degenerate with both water features and stellar contamination, a successful detection of a Rayleigh scattering slope could confirm the presence of a haze, providing an additional path towards breaking the opaque pressure level/metallicity degeneracy.
@@
VII Conclusions
We have carried out independent reductions of the first seven targets from the COMPASS program using the ExoTiC-JEDI code. We introduce a systematics model that accounts for variations in the trace morphology via the principal components of the relative pixel flux timeseries. We demonstrate that this systematics model reduces red noise in the residuals of the white light curves and improves the Akaike and Bayesian information criteria for low group-number observations. An analysis of the posteriors from the white light curve fitting confirms that the systematic effects are largest for the three-group observations. We also observe that the NRS1 detector suffers from systematics to a greater degree than NRS2.
@@ VII Conclusions
We derived new limits on metallicity and opaque pressure level assuming thermochemical equilibrium for each of the seven targets. We showed that the lower limits on metallicity are in broad agreement with those obtained in earlier studies with a few exceptions as noted in Section V.
@@
By comparing best-fit models from the set of small planets with JWST observations of transmission features in the near-infrared to hypothetical chemically-identical atmospheres around each of the COMPASS targets, we demonstrate that only the lowest mean molecular weight sub-Neptune or super-Earth atmospheres known, those of TOI-421 b and K2-18 b, would be detectable around all of the planets in the sample. In other words, we confirm that none of the planets in this sample have atmospheres that are analogs to K2-18 b or TOI-421 b. Notably this includes TOI-836 c, which is a close analog of TOI-421 b in mass and radius, though it is substantially cooler with an equilibrium temperature of compared to TOI-421 b’s . The atmosphere of TOI-421 b is ruled out for TOI-836 c at . If these planets have similar atmospheres, it would therefore imply the presence of high-altitude aerosols in TOI-836 c’s atmosphere similar to those observed for GJ 1214 b. With two to five transits of TOI-836 c observed with NIRISS/SOSS, a 50% steam/50% H2 atmosphere could be ruled out, while transits would be required to rule out a 100 Solar metallicity atmosphere with an opaque pressure level bar with NIRSpec/G395H. Using the same methods described in Section VI.2 we find that an atmosphere analogous to that of GJ-1214 b for TOI-836 c would be undetectable with fewer than several hundred NIRSpec/G395H transits due to its smaller transmission spectroscopy metric (Kempton et al., 2018). As a result, breaking the aerosol-metallicity degeneracy by detecting aerosols in the astmosphere of TOI-836 c would likely require the use of NIRISS/SOSS to detect the Rayleigh scattering slope at wavelengths shorter than those accessible to G395H.
The heavier atmospheres of TOI-270 d, L 98-59 b, and GJ 9827 d can all be ruled out for some of the seven COMPASS targets but are compatible with others, while the very high mean molecular weight, hazy atmosphere of GJ 1214 b is compatible with all seven of our spectra. In particular we note that TOI-270 d’s atmospheric composition is strongly ruled out for four of the seven targets (), weakly ruled out for TOI-836 c and TOI-776 b () and consistent only with TOI-776 c. Teske et al. (2025) also finds that TOI-776 c’s spectrum is compatible with TOI-270 d’s atmospheric composition, indicating the need for additional observations to rule out the hypothesis that these two planets, which have similar masses, radii, and equilibrium temperatures, have analogous atmospheres.
@@
In addition to the software cited in the text, we acknowledge the use of the following software in this work: astropy (Astropy Collaboration et al., 2022), scipy (Virtanen et al., 2020), numpy (Oliphant, 2007), emcee (Foreman-Mackey et al., 2019), matplotlib (Hunter, 2007), scikit-learn (Pedregosa et al., 2011), batman (Kreidberg, 2015).
Author Contributions
Co-Author contributions are as follows: TAG carried out the data analysis and led the writeup. NMB advised TAG throughout the work. NEB provided independent lower limits for the atmospheric metallicity and contributed text. All co-authors provided feedback on paper drafts and/or gave feedback during team meetings.